OpenStudy (anonymous):

hi please help me if we have tan(x)=-2 so what is sin(x),cos(x) and cot(x) ?

OpenStudy (anonymous):

sinx=-2 cox =1 cotx = 1/-2

OpenStudy (kira_yamato):

If tan(x) = -2, we can immediately say cot(x) = -1/2 There are 2 values for sin(x) and 2 corresponding values for cos(x), depending on your quadrant (which is determined by the range). Case I: 2nd Quadrant  By Pythagorean Theorem, hypotenuse ^ 2 = (-2)^2 + 1^2 hypotenuse = sqrt(5) sin(x) = 1/sqrt(5) and cos(x) = -2/sqrt(5) Case II: 4th quadrant  sin(x) = -1/sqrt(5) and cos(x) = 2/sqrt(5)